OpenStudy (anonymous):

WILL FAN AND MEDEL PLEASE HELP Select the functions that have identical graphs. 1) y=sin(3x+pi/6) 2) y=cos(3x-pi/6) 3) y=cos(3x-pi/3)

OpenStudy (gorv):

expand these functions and you will come to know

OpenStudy (gorv):

for \[\sin(A+B)=sinA*cosB+sinB*cosA\]

OpenStudy (gorv):

\[\cos(A-B)=cosA*cosB+sinA*sinB\] \[\cos(A+B)=cosA*cosB-sinA*sinB\]

OpenStudy (gorv):

use these 3

OpenStudy (anonymous):

sorry i dont stay online because i had a class to go to but thank you for helping